OpenStudy (anonymous):

f(x) =(x/8)-3, g(x)=x^3 (f*g)^-1=

OpenStudy (anonymous):

is that composite functions? \[(fog)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

pretty much substitute g(x) into the x of f(x)

OpenStudy (anonymous):

f(g(x))

OpenStudy (anonymous):

yeah but i got (x^(1/3)/8)-3

OpenStudy (anonymous):

is \[f(x)=\frac{ x }{ 8 }-3\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(fog)=\frac{ x^3 }{ 8 }-3\]

OpenStudy (anonymous):

just plug x ==> x^3

OpenStudy (anonymous):

it said it was wrong

OpenStudy (anonymous):

@chris00

OpenStudy (anonymous):

pic me the question

OpenStudy (anonymous):

ohh inverse

OpenStudy (anonymous):

\[y=f(x)=\frac{ x }{ 8 }-3\] y+3=x/8 8y+24=x then let x=y 8x+24=y=\[f(x)^{-1}\]

OpenStudy (anonymous):

so \[f(x)^{-1}=8x+24\]

OpenStudy (anonymous):

do the same for y=x^3

OpenStudy (anonymous):

that would be x^1/3

OpenStudy (anonymous):

y^(1/3)=x therefore \[g(x)^{-1}=x ^{1/3}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now\[(fog)^{-1}=8x ^{1/3}+24\]

OpenStudy (anonymous):

it says its wrong

OpenStudy (anonymous):

lol werid

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

( f o g)–1(x) = (g^–1 o f^ –1)(x)

OpenStudy (anonymous):

so we sub f(x) into the x function of g

OpenStudy (anonymous):

\[(fog)^{-1}=(8x+24)^{1/3}\]

OpenStudy (anonymous):

try that

OpenStudy (anonymous):

ah youre amazing!!!! thank you!!!!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

just try and know your properties. it a bit confusing http://www.purplemath.com/modules/fcncomp6.htm Thats a good link

OpenStudy (anonymous):

will do, thanks!

OpenStudy (anonymous):

: )